1.     A particular brand of petrol was used in 80 randomly chosen cars of the same model and age. The petrol consumption, x miles per gallon, was obtained for each car. The results are summarized by

∑x=1896 and ∑x2=45959.

Calculate a 99% confidence interval for the mean petrol consumption of all cars of this model and age.

 


1
Expert's answer
2022-04-21T04:15:11-0400

Since the population value of σ is unknown, we'll use the confidence interval for σ unknown.

Thus the 1 − α confidence interval is

"\bar{x}-t_{\alpha/2}\cfrac{s}{\sqrt{n}}<\mu<\bar{x}+t_{\alpha/2}\cfrac{s}{\sqrt{n}}"

where "t_{\alpha/2}" is the value from the t-distribution table for

"P(T>t_{\alpha/2})=\alpha/2."

For a 99% confidence interval we have α = 0.01 and α/2 = 0.005.

Using the table for df = n - 1 = 80 - 1 = 79 (degrees of freedom, n = 80 - the sample size):

"t_{0.005;79}=2.639."

The sample mean:

"\bar{x}=\cfrac{\sum x}{n}=\cfrac{1896}{80}=23.7."

The sample variance:

"s^2=\cfrac{\sum x^2-\cfrac{(\sum x)^2}{n}}{n-1}=\\
=\cfrac{45959-\cfrac{1896^2}{80}}{80-1}=12.96."

The sample standard deviation:

"s=\sqrt{12.96}=3.60."

Thus the 99% CI is

"23.7-2.639\cdot\cfrac{3.60}{\sqrt{80}}<\mu<23.7+2.639\cdot\cfrac{3.60}{\sqrt{80}}\\
22.64<\mu<24.76."